OpenStudy (secret-ninja):

medals if you give me a direct answer asap! The numbers 20, 21, 29. form a Pythagorean triple. Which other set of numbers is a Pythagorean triple? 21, 22, 30 40, 42, 58 60, 66, 87 70, 84, 116

OpenStudy (anonymous):

40, 42,58

OpenStudy (secret-ninja):

thank u!

OpenStudy (anonymous):

\[20^2+21^2=29^2\] \[40^2+42^2=58^2\]

OpenStudy (anonymous):

in case you were wondering how to do it

OpenStudy (secret-ninja):

thanks

OpenStudy (anonymous):

you find the set of numbers that the square of the largest number equals the squares of the 2 smaller numbers added together, none of the other number sets matched